# Phase-Stable Optical Fiber Links for Quantum Network Protocols

### v1

@@ Abstract
We demonstrate the distribution of single-photon-level pulses from a mode-locked laser source over a phase-stable fiber link, achieving an optical timing jitter of less than 100 as over 10 minutes of data accumulation. This stability enables a fidelity greater than 0.998 between two stabilized 2.1 km long deployed fiber links. Building on time and frequency metrology techniques traditionally used for high-stability optical atomic clock signal distribution, we use time and frequency multiplexing to achieve an isolation of quantum and classical channels by better than $10^{10}$. Our results mark a necessary step towards scalable, high-rate quantum networks with a provable quantum advantage.
@@ I Introduction
The future success of many quantum technologies that span computing, communications, sensing, and metrology will hinge on the ability to reliably distribute quantum entanglement over large-scale networks. Such distribution is not only central to practical applications like quantum key distribution (QKD) and distributed quantum computing, but also to foundational tests of quantum mechanics, including loophole-free Bell tests and demonstrations of quantum nonlocality over unprecedented distances. These tests are more than academic because they underpin the security of device-independent quantum protocols, where a violation of a Bell’s inequality certifies the presence of entanglement without trusting the devices themselves.
Optical fibers are natural conduits for all-photonic quantum networks due to their low-loss transmission and deployment flexibility. However, they are also highly sensitive to environmental perturbations, which can scramble photon polarization, alter arrival times, and shift optical phase, effects that degrade entanglement fidelity. In Bell-type experiments and quantum networking protocols, such disturbances reduce interference visibility and can even erase the nonclassical correlations that give these systems their advantage.
To address these challenges, we draw on techniques from precision time and frequency metrology, developed for comparing the world’s most accurate atomic clocks, and adapt them to stabilize optical paths for quantum state distribution. This approach directly supports protocols that rely on single-photon interference, including quantum-enhanced very long baseline interferometry [1, 2] and secure communication protocols such as twin-field quantum key distribution [3, 4], and certain quantum computing schemes [5]. Because interference is a fundamental property of quantum systems, phase stabilization forms a key base technology needed for many future quantum systems and applications.
The first quantum repeater architecture, proposed by Duan, Lukin, Cirac, and Zoller (DLCZ) [6], relies on quantum interference between the possible emissions of a single photon from two distant atomic ensembles. The presence or absence of a detected photon creates an entangled state between the two ensembles, but only if the photonic paths are indistinguishable and maintain a high level of phase stability. Any distinguishing information or optical path-length variation on the scale of tens of nanometers can degrade interference visibility and destroy entanglement.
In our quantum network in Boulder, CO, we aim to isolate and exploit the photonic component of the DLCZ scheme—specifically, the superposition of $|0\rangle$ and $|1\rangle$ number states generated with a beam splitter—to distribute entanglement over fiber. In this encoding, the photon’s spatial mode (its path) carries the quantum information. A key advantage is the scaling: the entangling rate of encodings that rely of entanglement between two photons, such as polarization or frequency, typically degrade linearly with transmission efficiency ($\eta$), whereas path entanglement can retain a $\sqrt{\eta}$ rate scaling, providing a more favorable loss tolerance as networks increase in length and number of nodes [7]. The trade-off is that path entanglement imposes extremely stringent requirements on optical path-length stability, with fluctuations of only tens of nanometers sufficient to wash out interference and reduce fidelity.
In this demonstration, we achieve three critical performance benchmarks for path-entangled quantum networking over a deployed fiber link: (1) optical timing jitter below 100 as over 10 minutes, (2) indistinguishability between fiber paths exceeding 99.6%, and (3) isolation between classical and quantum channels greater than $8\times 10^{10}$ (ratio of number of photons in the classical channel to the number of photons that couple from the classical channel to the quantum channel). These capabilities pave the way for high-rate, high-fidelity entanglement distribution over dedicated municipal networks needed for next-generation distributed quantum applications.
II Path Entanglement
The DLCZ scheme for quantum networking utilizes entanglement between the possible physical paths a photon traverses, leading to the term path entanglement. In this work we consider path entanglement that is generated by a single photon interacting with a 50/50 beam splitter. The photon can exit the beam splitter into either path $A$ or path $B$, leading to the maximally-entangled Bell state:
@@ II Path Entanglement
$|\psi_{\text{path}}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle_{A}|1\rangle_{B}+e^{i\Delta\phi}|1\rangle_{A}|0\rangle_{B}\right).$ (1)
@@
The phase term, $\Delta\phi$, represents the difference in accumulated optical phase between the two paths with a length difference of $\Delta L$. The relationship between phase and length, $\Delta\phi=k\times\Delta L$, where $k$ indicates the wavevector, demonstrates the sensitivity of the quantum state to perturbations in optical path length. Therefore, to distribute path entanglement over optical fiber links with high fidelity, it is necessary to actively stabilize the fiber length.
Entanglement swapping can be used to entangle the relative paths of two remote nodes as shown in Fig. 1. Nodes A and B each use a single-photon source to produce a path-entangled state. One path is kept locally and used for quantum protocols and the other path is sent to a quantum repeater, which overcomes entanglement fidelity degradation due to fiber network loss by heralding entanglement between nodes. That is, when a single photon is detected, it is impossible to know which source it came from. This entangles the path states of sources A and B. A counter-propagating classical light source located at the entanglement swapping beamsplitter is used to actively stabilize both fiber paths. The classical laser also distributes a common optical phase reference across the network and synchronizes remote single-photon source events. The classical communication of projective measurements is not shown, though it is necessary to verify entanglement and implement quantum protocols.
Figure 1: A two-node path-entangled quantum network with a classical channel (purple lines) and quantum channel (red lines). A classical laser is used to stabilize network fiber links and to synchronize single-photon events at each node. A quantum repeater helps overcome network link losses and heralds entanglement. Entanglement heralding and local state projective measurements are recorded locally and transmitted over a different classical communication line (not shown).
@@
To achieve high-fidelity (> 0.99) path-state entanglement distribution, several conditions must be met. Phase stability is crucial, as even small fluctuations in optical path length can introduce phase noise that limits the measurement fidelity. A phase change of $\Delta\phi=0.1$ radians reduces the fidelity to about 0.9975 (discussed in more detail in section III.2). Additionally, the optical paths from source to measurement must be highly indistinguishable, that is, there must be no identifying information about which source a photon came from that reduces the interference visibility at the repeater. Photons from the two sources must have the same polarization, spectral and temporal profiles, and arrive at the same time within femtoseconds. Finally, a high degree of isolation between quantum and classical channels is required to ensure that the relatively high-power stabilization laser does not overwhelm the single-photon detectors or lower the fidelity with noise photons.
Implementing a full entanglement swapping demonstration is beyond the scope of this manuscript. Instead, we focus on independently stabilizing the the two deployed fiber paths that would connect nodes A and B to the swapping station in a Mach-Zehnder interferometer as shown in Fig. 2.
@@ III The classical channel
Figure 2: Experimental setup for the co-propagation of a quantum single-photon channel and a classical single-frequency phase-stabilization channel. The channels are isolated from one another by time-interleaving the classical and quantum channels using a mechanical chopper. The quantum signal, generated by attenuating a 500 MHz repetition rate, 1550 nm mode-locked laser to an average of <1<1 photon per pulse, is combined with the classical channel via a 50/50 optical beamsplitter. Length compensation is achieved using a fiber stretcher and motorized delay line. Two independently stabilized fiber links are employed to assess the performance of stabilization on both channels by measuring residual phase noise power spectral density on the classical channel and optical fringe contrast on the quantum channel. SNSPD - superconducting nanowire single-photon detector, AOM - acousto-optic modulator, PID - proportional/integral/derivative loop filter
@@ III.1 Optical Phase Stabilization
The classical channel consists of a narrow-linewidth continuous wave (CW) laser at 1542 nm, stabilized to an optical reference cavity that enables a full-width half-maximum linewidth, $\Delta\nu$, less than 10 mHz [10]. The choice of stabilization laser is important since the coherence time of the laser, $\tau_{\text{coh}}=1/(\pi\Delta\nu)$, must far exceed the transit time of light through the optical fiber. If the phase of the laser changes significantly over the fiber propagation time, it is impossible to distinguish between laser and fiber noise, and the laser noise is written onto the fiber.
@@
The CW light is chopped by a 6 kHz chopper with 50% duty cycle, which is anti-synchronized with a 25% duty cycle chopper located before the single photon detector in the quantum channel. After chopping, a fraction of the CW light is split off by a beam splitter and used as a stable interferometric reference, while the remainder of the light traverses the optical fiber. At the remote end, about 4% (Fresnel reflection) of the signal is retro-reflected from the flat fiber tip such that it returns through the fiber and interferes with the reference light at the local end, generating an error signal that is used to feedback and suppress the fiber noise. An acousto-optic modulator (AOM) is located at the remote end of the fiber, which shifts the reflected optical frequency by twice the RF drive frequency (80 MHz). The AOM is advantageous because it shifts the interference signal frequency away from DC, which avoids DC noise sources such as photodetector flicker noise and renders the stabilization immune to laser power fluctuations. Additionally, locating the AOM at the end of the link helps to distinguish reflections off the end of the fiber from unwanted reflections of other locations along the link, further increasing the signal to noise ratio [11].
@@
A fiber stretcher acts as the fast feedback actuator (> 20 kHz bandwidth) and a motorized delay line acts as a secondary actuator for slow and large length changes. It is important to use an actuator that affects the optical path length and not only the optical phase (e.g., an acousto-optic modulator). This is because we must simultaneously control both the optical phase and also the timing of the propagating wavepackets. A stretcher and free space delay line also allows us to stabilize the fiber at one wavelength and expect a very high level of stabilization for a different wavelength.
@@
Time multiplexing of the classical and quantum channels adds complexity to the feedback electronics. To prevent unlocks while the chopper blocks the light, we use an FPGA-based feedback algorithm [12] where we implement a digital lowpass filter to limit the response bandwidth to less than the chopper speed.
@@
Optical phase distribution across the network: For future quantum network experiments, the remaining 96% of the classical light that is not reflected from the remote end of the fiber for stabilization may be used to distribute a common optical phase across all nodes [7, 13]. Such a phase reference may be used to lock an optical frequency comb at each node, used to pump the generation of single-photon states. The comb may also be used to generate a radio frequency signal to clock electronic circuits that is coherent with the distributed optical signal.
III.2 Fiber link stability as measured by a classical signal
@@
Figure 3: (a) Phase noise power spectral density (PSD) of the 1550 nm optical beatnote between two stabilized optical fiber links (2.1 km and 2 km), showing the frequency-dependent phase fluctuations as a function of chopping rate. (b) Corresponding integrated root-mean-square (RMS) phase deviation, illustrating the cumulative RMS phase noise as a function of integration time.
@@
$\Delta\phi_{rms}=\sqrt{\int_{1/50\mu s}^{1/\tau_{int}}S_{\phi}(f)df}.$ (2)
To measure $S_{\phi}(f)$, we independently stabilize two fiber links, one with a length of about 2 km and the other about 2 m, with the same 1542 nm chopped CW laser. We assume that the longer fiber contributes nearly all of the added phase noise. To classically test the phase noise of the quantum channel, we split and transmit optical pulses derived from a stabilized mode-locked laser through each fiber and interfere the outputs on a high-speed photodetector (1.5 GHz bandwidth) at the fiber outputs. The mode-locked laser is based on a 500 MHz repetition rate Er/Yb:glass oscillator [14], whose carrier-envelope offset frequency is stabilized [15, 16, 17] and whose spectrum is phase-stabilized to the 1542 nm CW laser. The pulse is filtered by a fiber Bragg grating centered at 1550 nm and with a full-width half-maximum bandwidth of 1 nm.
Figure 3 shows the measured phase noise power spectral density, and corresponding integrated phase noise for different chopper rates of the 1550 nm quantum channel. A slow chopper speed limits the bandwidth of the fiber stretcher feedback and will therefore be inadequate to fully suppress high-bandwidth noise (100 Hz to 2 kHz) that the fiber encodes on the light. This is seen in the 4 kHz phase noise trace (blue) as peaks near 1 kHz, which contribute much of the integrated phase noise. Conversely, a high chopping speed introduces additional noise across all frequencies due to the mechanical timing jitter of the chopping wheel, which increases at higher speeds. This is evident in the red and green phase noise spectra as chopping increases from 4 kHz to 10 kHz, and the effect is absent when the CW laser is not chopped.
@@
$F(\Delta\phi_{rms})=\left|\langle\psi_{\text{path}}|\psi_{\text{path}}^{\prime}\rangle\right|^{2}=\frac{1}{2}\left(1+\cos(\Delta\phi_{rms})\right),$ (3)
where $|\psi_{\text{path}}\rangle$ is the state without phase noise and $|\psi_{\text{path}}^{\prime}\rangle$ is the state with added phase noise. The fidelity is a measure of the indistinguishability of two states produced when one photon propagates along two paths, according to Eq. 1. The relative phase change over some amount of time between two optical paths degrades the fidelity.
The unchopped experiment achieves an optical timing jitter well below 50 as, corresponding to $F>0.999$. Among the chopped cases, 6 kHz gives the best results with an optical timing jitter of about 70 as, corresponding to $F\approx 0.998$, out to an integration time of at least 20 seconds. Although this is evidence that optical chopping degrades the fidelity, it is required for high classical/quantum channel isolation in this demonstration.
The integration time defines how long we can assume a nearly fixed phase relationship between fiber paths, setting the maximum duration over which we can perform projective measurements reliably. To achieve higher state fidelity, measurements can be taken over shorter time windows, with phase recalibration between windows. This approach allows multiple measurements to be cascaded for longer protocols or measurements that require additional averaging.

### v2

@@ Abstract
We demonstrate the distribution of single-photon-level pulses from a mode-locked laser source over a phase-stable fiber link, achieving an optical timing jitter of less than 100 as over 10 minutes of data accumulation. This stability enables a fidelity greater than 0.998 between two stabilized 2.1 km long deployed fiber links. Building on time and frequency metrology techniques traditionally used for high-stability optical atomic clock signal distribution, we use time and frequency multiplexing to achieve an isolation of quantum and classical channels of $8\times 10^{10}$. Our results mark a necessary step towards scalable, high-rate quantum networks with a provable quantum advantage.
@@ I Introduction
The future success of many quantum technologies, including computing [15], communications [12, 21, 40], sensing, and metrology [13, 8], hinges on the ability to reliably distribute quantum entanglement across a network. Because quantum states cannot be copied, transmission loss poses a fundamental challenge for long-distance quantum communication. Overcoming this limitation requires quantum repeaters, which enable efficient distribution of entanglement over lossy channels. The first quantum repeater architecture, proposed by Duan, Lukin, Cirac, and Zoller (DLCZ) [9], relies on quantum interference between two indistinguishable pathways corresponding to the emission of a single photon from spatially separated atomic ensembles. After interfering these two paths, the detection of a single photon heralds the creation of an entangled state between the ensembles (as shown in Fig. 1 a)), provided that the photonic paths are indistinguishable. A central challenge for DLCZ-based architectures is their sensitivity to optical phase fluctuations. Because the distributed entangled state depends critically on the relative phase between the photonic paths, all fiber links in the network effectively form a large interferometer. Path-length fluctuations on the order of tens of nanometers are sufficient to wash out interference and significantly reduce entanglement fidelity. Entanglement purification [10, 3, 30] can be used to create higher fidelity entanglement from multiple lower fidelity entangled states. However, the number of required copies increases as the initial entangled state fidelity drops. Consequently, even modest reductions in phase-noise-induced infidelity can substantially enhance repeater throughput and link distance.
In this work we demonstrate a telecom-wavelength, multi-fiber, quantum-compatible phase-stabilization technique for DLCZ-style quantum repeaters that performs nearly an order of magnitude better in interference visibility compared to previously reported approaches. Using methods adapted from the atomic clock and optical time-transfer communities, we independently stabilize two 2.1 km single-mode fibers in the interferometric configuration required for entanglement swapping in a quantum repeater. Excess noise was intentionally introduced by aerially suspending the fibers, in contrast to equivalent underground deployments, providing a stringent test of stabilization performance under realistic environmental perturbations. In this deployed-fiber demonstration, we achieve three critical performance benchmarks for path-entangled quantum networking: (1) optical RMS timing jitter around 70 as over 20 seconds (corresponding to a global phase noise of $\Delta\phi_{\mathrm{RMS}}=0.085$ radians) and a timing drift of less than 100 as over 10 minutes, (2) path indistinguishability exceeding 99.6 $\%$, and (3) isolation between classical stabilization signals and the quantum channel greater than $8\times 10^{10}$, defined as the ratio of photons in the classical channel to those coupling into the quantum channel. Our approach is suitable for both matter-based and high event rate all-optical DLCZ repeater architectures needed to support municipal-scale distributed quantum applications.
A key advantage of the DLCZ architecture lies in its scaling behavior: repeater protocols based on two-photon encodings (e.g., polarization or frequency) typically exhibit entanglement rates that scale linearly with the transmission efficiency, $\eta$, whereas DLCZ-type protocols retain a $\sqrt{\eta}$ scaling[33]. This more favorable scaling makes DLCZ-style repeaters particularly attractive as network length and complexity increase. As a result, state-of-the-art quantum repeater demonstrations [20, 34] are based on variants of the DLCZ protocol. However, simultaneously stabilizing multiple deployed fibers is challenging and typically requires using bright classical light to measure relative path length changes. However, this bright light must coexist with the single-photons used to perform the entanglement swapping. To prevent unwanted crosstalk between the classical and quantum channels, a high degree of isolation is required.
Fluctuations or drifts in the relative phase between different paths place a fundamental upper bound on the achievable entanglement fidelity, $F$. For a path-entangled state, this bound is given by
@@
$F(\Delta\phi_{\mathrm{RMS}})=\left|\langle\psi_{\text{path}}|\psi_{\text{path}}^{\prime}\rangle\right|^{2}=\frac{1}{2}\left(1+\cos(\Delta\phi_{\mathrm{RMS}})\right),$ (1)
where $\Delta\phi_{\mathrm{RMS}}$ is the root-mean-square phase error. Recent experimental demonstrations highlight the impact of this limitation. In ref. [20], the repeater links exhibited an RMS phase error of 0.297 radians, corresponding to a maximum achievable fidelity of $\sim 0.98$. Similarly, ref. [33] reported an RMS phase error of 0.609 radians, limiting the fidelity to $\sim 0.91$. The stabilization requirements for a phase-sensitive quantum repeater are also subtly different than those needed for other quantum communication protocols, such as twin-field quantum key distribution (TF-QKD)[21, 40]. In recent TF-QKD demonstrations the light in the quantum channel is bright enough to use to directly track the phase, removing the need for a separate stabilization light source. Additionally, the needed corrections can be done via post processing. In contrast, for general purpose DLCZ-style repeaters phase fluctuations must be dynamically corrected either by dynamically adjusting the path lengths involved or by applying corrections to the stored qubits at the end stations.
II All-optical DLCZ architecture
Although most quantum repeater architectures rely on matter-based qubits at the end stations, an all-optical implementation of the DLCZ repeater is possible. In this approach, matter-based qubits are replaced by on-demand single-photon sources and linear optical elements, with entanglement encoded in the photonic path degrees of freedom. In principle, single-photon sources can operate at gigahertz rates, leading to substantially higher entanglement swapping event rates. Available all-optical quantum memories[17] mean that this architecture is best suited for high-performance metropolitan-scale quantum networks. The techniques developed in this work show that optical paths in a high-bandwidth, all-optical DLCZ-based quantum repeater can be stabilized with high visibility and strong isolation between classical and quantum channels. This represents a critical first step toward building scalable, high-performance, all-photonic entanglement swapping in metropolitan-scale networks.
[…]
$|\psi_{\text{path}}\rangle=\frac{1}{\sqrt{2}}\left(|0\rangle_{a}|1\rangle_{b}\pm e^{i\Delta\phi}|1\rangle_{a}|0\rangle_{b}\right).$ (2)
@@ II All-optical DLCZ architecture
The phase term, $\Delta\phi$, represents the difference in accumulated optical phase between the two paths $a^{\prime}$ and $b^{\prime}$ with a length difference of $\Delta L$. The relationship between phase and length, $\Delta\phi=k\times\Delta L$, where $k$ indicates the wavevector, demonstrates the sensitivity of the quantum state to perturbations in optical path length. Therefore, to distribute path entanglement over optical fiber links with high fidelity, it is necessary to actively stabilize the fiber length. An additional overall $\pi$ phase shift is introduced depending on which detector in the BSM fires. Using feed forward, this extra phase shift can be corrected.
The two remote nodes can in principle couple each of their entangled paths to an atomic ensemble to create entanglement between their respective ground and excited states. Alternatively, the remote nodes can either store their quantum states in a phase-stable variable optical delay that acts as a memory, or they can directly measure their path entangled qubit using a weak local oscillator as shown in Figure 1 b). A weak local oscillator and single-photon detector can approximate single-qubit measurements in a path-entangled encoding[25, 1]. A weak local oscillator, for instance, can perform the projective measurements needed to show a violation of a Bell inequality needed for advanced device-independent quantum communication protocols as shown in [1, 2, 38]. The benefit of an all-optical DLCZ repeater is that the heralded photon source may operate at trial rates orders of magnitude higher than comparable matter-based systems, but comes with the additional requirement that the local oscillators at each location share the same phase.
@@
To achieve high-fidelity (> 0.99) path-state entanglement distribution, several conditions must be met. Phase stability is crucial, as even small fluctuations in optical path length can introduce phase noise that limits the measurement fidelity. A phase change of $\Delta\phi=0.1$ radians reduces the fidelity to about 0.9975 (discussed in more detail in section III.2). Additionally, the optical paths from source to measurement must be highly indistinguishable, that is, there must be no identifying information about which source a photon came from that reduces the interference visibility at the repeater. Photons from the two sources must have the same polarization, spectral and temporal profiles, and arrive at the same time within femtoseconds. Finally, a high degree of isolation between quantum and classical channels is required to ensure that the relatively high-power stabilization laser does not overwhelm the single-photon detectors or lower the fidelity with noise photons. Implementing a full entanglement swapping demonstration is beyond the scope of this manuscript. Instead, we focus on independently stabilizing the two deployed fiber paths that would connect nodes A and B to the swapping station in a Mach-Zehnder interferometer (as shown in Fig. 2) that would enable such a demonstration. We use an attenuated laser at 1550 nm to simulate a heralded telecom photon source operating with a bandwidth of 1 nm at a repetition rate of 500 MHz and average photon number of 0.08. A major benefit of our scheme is that the stabilization signal that is distributed can in principle be used to lock independent weak local oscillators needed to carry out the measurements on the path-entangled qubits.
@@ III The classical channel
Figure 2: Experimental setup used for the phase stabilization and noise measurement of a deployed 2.1 km fiber link. This setup operates by counter-propagating light from the single-photon quantum channel, in the same spatial mode as classical single-frequency light from the phase-stabilization channel. The channels are isolated from one another by time-multiplexing the classical and quantum channels using a mechanical chopper. The quantum signal, generated by attenuating a 500 MHz repetition rate, 1550 nm mode-locked laser to an average of <1<1 photon per pulse, is combined with the classical channel via a 50/50 optical beamsplitter. Length compensation is achieved using round trip measurement of the fiber phase noise, used in a feedback loop, and actuated via a fiber stretcher and motorized delay line. Two independently stabilized 2.1 km fiber links are employed to assess the stabilized performance of both channels by measuring the phase noise power spectral density of the quantum channel. A third short link is used to connect the output of the free space interferometer with a superconducting nanowire single-photon detector (SNSPD) to assess the fringe visibility of single photons traversing the stabilized fiber link. In the figure, AOM - acousto-optic modulator, PID - proportional/integral/derivative loop filter.
@@ III.1 Optical Phase Stabilization
The classical channel consists of a narrow-linewidth continuous wave (CW) laser at 1542 nm, stabilized to an optical reference cavity that enables a full-width half-maximum linewidth, $\Delta\nu$, less than 10 mHz [23]. The choice of stabilization laser is important since the coherence time of the laser, $\tau_{\text{coh}}=1/(\pi\Delta\nu)$, must far exceed the transit time of light through the optical fiber. If the laser phase drifts over the fiber transit time, the system cannot separate laser noise from fiber noise, causing the laser’s phase fluctuations to be imprinted onto the link instability.
@@
The CW light is chopped by a 6 kHz chopper with 50% duty cycle, which is anti-synchronized with a 25% duty cycle chopper located before the single photon detector in the quantum channel. We use mechanical choppers because they achieve theoretically perfect extinction while other modulators (e.g., EOMs or AOMs) can achieve only up to 60 dB extinction under very carefully controlled conditions. After chopping, a fraction of the CW light is split off by a beam splitter and used as a stable interferometric reference, while the remainder of the light traverses the optical fiber. At the remote end, about 4% (Fresnel reflection) of the signal is retro-reflected from the flat fiber tip such that it returns through the fiber and interferes with the reference light at the local end, generating an error signal that is used to feedback and suppress the fiber noise. An acousto-optic modulator (AOM) is located at the remote end of the fiber, which shifts the reflected optical frequency by twice the RF drive frequency (80 MHz). The AOM shifts the interference signal frequency away from DC, which avoids DC noise sources such as photodetector flicker noise and renders the stabilization immune to laser power fluctuations. Additionally, placing the AOM at the end of the link frequency-shifts the returning light, allowing reflections from the fiber end to be distinguished from unwanted reflections within the link. This separation improves the signal-to-noise ratio of the stabilization signal [39].
@@
A fiber stretcher serves as the fast feedback actuator (>20 kHz bandwidth), while a motorized delay line provides slow, large-range path-length adjustments. It is essential that the actuator modify the optical path length rather than only the optical phase (as would an acousto-optic modulator), because the stabilization must control both the phase and the arrival time of the propagating wavepackets. Using a fiber stretcher in conjunction with a free-space delay line also allows the link to be stabilized at one wavelength while achieving similarly high levels of stabilization for another.
@@
Time multiplexing of the classical and quantum channels adds complexity to the feedback electronics. To prevent the servo from unlocking when the chopper blocks the stabilization light, we use an FPGA-based feedback algorithm [26] where we implement a digital low-pass filter to limit the response bandwidth to less than the chopper speed. This ensures that the servo responds only to the averaged phase error over multiple chopper cycles, rather than to the brief signal dropouts.
@@
Optical phase distribution across the network: For future quantum network experiments, the remaining 96% of the classical light that is not reflected from the remote end of the fiber for stabilization may be used to distribute a common optical phase across all nodes [33, 6]. Such a phase reference may be used to lock an optical frequency comb [11] at each node, used to pump the generation of single-photon states. The comb may also be used to generate a radio frequency signal to clock electronic circuits that is coherent with the distributed optical signal.
III.2 Quantum channel stability as measured by a classical signal
@@
Figure 3: (a) Phase noise power spectral density (PSD) of the 1550 nm optical beatnote between a 2.1 km and 2 m stabilized optical fiber links, showing the frequency-dependent phase fluctuations as a function of chopping rate. The light outputs from the two fibers are heterodyned to enable the combined measurement noise of both links, with the 2 m fiber contributing much less. The gray shaded area shows the noise floor when the two fibers are matched in length. (b) Corresponding integrated root-mean-square (RMS) phase deviation, illustrating the cumulative RMS phase noise as a function of integration time.
@@
$\Delta\phi_{\mathrm{RMS}}=\sqrt{\int_{1/50\mu s}^{1/\tau_{int}}S_{\phi}(f)df}.$ (3)
To measure $S_{\phi}(f)$, we independently stabilize two fiber links, one with a length of 2.1 km and the other about 2 m, with the same 1542 nm chopped CW laser. We assume that the longer fiber contributes nearly all of the added phase noise. To classically test the phase noise of the quantum channel, we split and transmit optical pulses derived from a stabilized mode-locked laser through each fiber and interfere the outputs on a high-speed photodetector (1.5 GHz bandwidth) at the fiber outputs. The mode-locked laser is based on a 500 MHz repetition rate Er/Yb:glass oscillator [24], whose carrier-envelope offset frequency is stabilized [35, 29, 16] and whose spectrum is phase-stabilized to the 1542 nm CW laser. The pulse is filtered by a fiber Bragg grating centered at 1550 nm and with a full-width half-maximum bandwidth of 1 nm.
The goal of the present work is not to improve the performance of earlier fiber link demonstrations [22, 39, 6], but to provide a stabilized infrastructure with quantum-compatible performance for future path-entanglement and multi-node interference experiments over deployed networks. In contrast to classical time–frequency links that typically aim to maximize fiber length, our approach deliberately limits the link length to achieve a target quantum-state fidelity, as phase noise accumulated across the full Fourier spectrum directly leads to decoherence and reduced interference visibility of path-entangled states. The demonstrated architecture therefore addresses a complementary regime in which high phase stability and indistinguishability are prioritized over absolute distance, enabling scalable quantum networking protocols under realistic coexistence constraints.
In Figure 3 we perform classical measurements to characterize the quantum channel. The figure shows the measured phase noise power spectral density, and corresponding integrated phase noise for different chopper rates of the 1550 nm quantum channel, while stabilized with the 1542nm classical channel. A slow chopper speed limits the bandwidth of the fiber stretcher feedback and will therefore be inadequate to fully suppress high-bandwidth noise (100 Hz to 2 kHz) that the fiber encodes on the light. This is seen in the 4 kHz phase noise trace (blue) as peaks near 1 kHz, which contribute much of the integrated phase noise. Conversely, a high chopping speed introduces additional noise across all frequencies due to the mechanical timing jitter of the chopping wheel, which increases at higher speeds. This is evident in the red and green phase noise spectra as chopping increases from 4 kHz to 10 kHz, and the effect is absent when the CW laser is not chopped.
@@
The righthand axis shows the optical timing jitter associated with the integrated phase noise, $\Delta\tau_{\mathrm{RMS}}=\Delta\phi_{\mathrm{RMS}}/2\pi f_{0}$, where $f_{0}$ is the optical frequency, and theoretical path state fidelity is given in Eq. 1. The fidelity is a measure of the indistinguishability of two states produced when one photon propagates along two paths, according to Eq. 2. Temporal relative phase change between two optical paths can degrade the fidelity.
For fiber length stabilization, in the absence of chopping, we achieve an optical timing jitter well below 50 as, which theoretically maps to a fidelity of $F>0.999$. Among the different chopper rates, 6 kHz gives the best results with an optical timing jitter of about 70 as, corresponding to $F\approx 0.998$, out to an integration time of at least 20 seconds. Although this is evidence that optical chopping degrades the fidelity, chopping is required for high classical/quantum channel isolation in this demonstration.
The integration time defines how long we can assume a nearly fixed phase relationship between fiber paths, setting the maximum duration over which we can perform projective quantum measurements reliably. To achieve higher state fidelity, measurements can be taken over shorter time windows, with phase recalibration between windows. This approach allows multiple measurements to be cascaded for longer protocols or measurements that require additional averaging.
III.2.1 In-loop vs. out-of-loop phase noise analysis
A careful phase-noise analysis of any stabilized system requires a dedicated measurement channel that is not involved in the feedback loop. In this experiment, each fiber is independently stabilized based on the phase information seen on its in-loop photodetector. A separate out-of-loop detector, located at the opposite beam-splitter port from the SNSPD, is used to characterize the residual phase noise of the two fiber links together. Importantly, this out-of-loop photodetector is not used in any feedback loop and therefore provides an unbiased measure of the stabilized system performance.
In-loop measurements, such as those obtained from detectors used in the feedback loop, can underestimate the true residual phase noise. This is because feedback suppresses the measured error signal while remaining insensitive to noise sources that are written onto the fiber by the actuator, including technical noise that appears as a valid phase signal, such as detector noise or electrical pickup such as ground loops, as well as phase fluctuations that are not common between the classical stabilization channel and the quantum channel. By contrast, an out-of-loop measurement captures all phase noise that remains on the transmitted optical field, including contributions outside the control bandwidth and noise arising from non-common-mode optical paths.
A particularly stringent out-of-loop test is obtained by independently stabilizing two fiber links and interfering them at a common beam splitter, such that the measured phase noise reflects the combined residual instability of both links. This approach provides a conservative assessment of phase stability and is directly applicable to quantum networking scenarios in which independent stabilized links must maintain mutual phase coherence.